OpenStudy (anonymous):

Alyssa gave the following explanation for factoring the trinomial x2 – 3x – 4: “I need to find factors of –4 that add to –3. 4 and –1 multiplies to –4 and adds to –3 so my factors are (x + 4)(x – 1).” Part 1: Are Alyssa’s factors correct? If not, explain where she made her mistake. (2 points) Part 2: What could Alyssa have done to check the accuracy of her factors before submitting her Module 3 test? (2 points)

OpenStudy (maheshmeghwal9):

Alyssa isn't correct.

OpenStudy (maheshmeghwal9):

She is wrong here; "4 and –1 multiplies to –4 and adds to –3 so my factors are (x + 4)(x – 1).” How can it be \[4-1=-3.\]? It should be -4 & 1 such as\[\color{red}{-4+1=-3} {and} \color{blue}{-4*1=-4.}\]

OpenStudy (anonymous):

can you help with the 2nd part too

OpenStudy (maheshmeghwal9):

just multiply the factors & u will not get ur actual trinomial.

OpenStudy (maheshmeghwal9):

it is a checking method

OpenStudy (anonymous):

ohhh okay thank so much

OpenStudy (maheshmeghwal9):

yw:)